# Latent Space Imaging

### v1

@@ Abstract
Digital imaging systems have classically been based on brute-force measuring and processing of pixels organized on regular grids. The human visual system, on the other hand, performs a massive data reduction from the number of photo-receptors to the optic nerve, essentially encoding the image information into a low bandwidth latent space representation suitable for processing by the human brain. In this work, we propose to follow a similar approach for the development of artificial vision systems.
@@
Latent Space Imaging is a new paradigm that, through a combination of optics and software, directly encodes the image information into the semantically rich latent space of a generative model, thus substantially reducing bandwidth and memory requirements during the capture process.
@@
We demonstrate this new principle through an initial hardware prototype based on the single pixel camera. By designing an amplitude modulation scheme that encodes into the latent space of a generative model, we achieve compression ratios from 1:100 to 1:1,000 during the imaging process, illustrating the potential of latent space imaging for highly efficient imaging hardware, to enable future applications in high speed imaging, or task-specific cameras with substantially reduced hardware complexity.
keywords:
Computational Imaging, Generative Models, Single Pixel Imaging
1 Main
@@
Conventional imaging systems are based on the brute-force sampling of scene information on regular grids. Biological vision systems operate very differently: for example the human visual system has around 120 million rods and 7 million cones [1], but the information from these photoreceptors is distilled by retinal processing into neural signals that are transmitted over the optic nerve with only 0.7-1.7 million axions [2]. This compression forms an encoding of the image information into a latent space that is compatible with the neural architecture of the human brain.
@@
Here we propose Latent Space Imaging (LSI), a new paradigm that aims to bring this concept to the realm of cameras and artificial vision systems. Specifically, we propose to utilize the combination of linear optical encoding for drastic data reduction, paired with a small amount of non-linear computational processing to directly encode image information into the latent space of a generative model. Overall, this combination drastically reduced the amount of sensor data that needs to be captured relative to a conventional 2D image sensor.
@@
Latent Space Imaging builds and improves upon the recent trend for co-designing both optics and reconstruction algorithms. This joint optimization has yielded favorable outcomes in computational imaging and downstream computer vision tasks. These include color image restoration [3], microscopy [4, 5, 6], monocular depth imaging [7, 8, 9, 10], super-resolution, extended depth of field [11, 12], time-of-flight imaging [13], and high-dynamic range imaging [14], among others. The enhancements observed are attributed to the ability of such systems to optically encode relevant scene information into sensor data, with the computational post-processing acting as a decoder. The crucial innovation in LSI is that the code represents the rich latent space of a generative model, with that model acting as the decoder module.
While there are a number of possible ways to optically implement the LSI paradigm, our first prototype utilizes the single-pixel imaging (SPI) concept [15, 16, 17, 18]. SPI was originally developed under the Compressed Sensing (CS) framework [19], and indeed LSI and CS share the common goal of pre-sensing data reduction. However, where CS utilizes a generic (usually random) encoding of the image, LSI uses codes that are specifically designed to embed the image into the rich latent space of a generative model in the form of a StyleGAN [20, 21, 22]. StyleGANs are renowned for their ability to synthesize highly realistic face images, editing and image-to-image translation applications. This capability stems from their semantically rich, disentangled, well-behaved, and compact latent space, making them an ideal architecture for LSI.
To showcase our framework, we tackle the particularly challenging task of full image reconstruction of human faces with the aim of accurate identity retrieval. Considering the distinctiveness of facial features, which present considerable challenges for existing compressed imaging techniques state-of-the-art usually suffer from blurred features and an overall loss of detail. LSI on the other hand consistently produces detailed, aesthetically pleasing faces that remain true to the original signal. Our experiments on this particularly challenging task illustrate potential for even higher compression ratios in easier tasks such as object detection or classification. Overall, the simplicity of LSI offers vast potential for expansion across a broad spectrum of applications in computer vision.
2 Results
2.1 Latent Space Imaging
Refer to caption a b c e d
Figure 1: (a) Traditional imaging methods require a large number of measurements and sensors with extensive pixel counting capabilities. b) Compressed sensing techniques typically act as low-pass filters, and when the measurements are compressed below $1\%$ of the original pixel counting they fail to reconstruct highly detailed images, even with the aid of sophisticated deep neural networks during the reconstruction process. (c) By encoding the image signal into the latent space of a generative model, surprisingly high compression rates can be achieved, while maintaining distinct fine details. (d) A linear optical and a non-linear  digital encoder are jointly optimized to acquire compressed data, and encode it into the latent space provided by the inversion network $\mathcal{E}$ by minimizing the loss function $\mathcal{L}_{lat}$. During reconstruction, the compressed data is fed into the generative model $\mathcal{G}$ (StyleGAN2) in order to reconstruct the image. The lock symbol represents a fixed model that is not being optimized. (e) We illustrate one possible implementation of the Latent Space Imaging technique using a single-pixel framework. An objective lens focuses the image onto the Digital Micromirror Device (DMD), which is responsible for implementing the learned mask to spatially modulate the incoming signal. This is followed by a relay lens, which focuses the modulated signal onto a photodiode (SPD) responsible for integrating the signal. Utilizing time multiplexing, we can retrieve the necessary measurements.
@@ 2.1 Latent Space Imaging
Traditional imaging captures a regular $m\times n$ grid of measurements (Fig. 1a), while Compressed Sensing (CS) uses a collection of random linear codes to leverage the information redundancy inherent in real images (Fig. 1b) in combination with a reconstruction software module. In contrast, we demonstrate that meaningful information can be significantly more compressed in the latent space of generative models (Fig. 1c). Our method allows for direct imaging into this space, achieving high quality image reconstructions from only $1.0\%$ to $0.1\%$ of the measurements compared to the original pixel count. Unlike traditional approaches that learn to interpolate missing information, LSI enables detailed reconstruction focusing on the extraction of key domain features through a combined optical and digital encoding process.
@@
As depicted by Fig. 1d, our objective is to learn an optically encoded mapping
@@
where the input image $\textit{I}\in\mathbb{R}^{mn}$ is mapped to a tensor $\textbf{L}\in\mathbb{R}^{k\times l}$ representing the latent space of a generative model. This mapping is achieved through the combination of a co-designed (linear) optical encoder $\textbf{O}\in\mathbb{R}^{mn\times d}$ and a digital (nonlinear) encoder $\mathcal{D}_{\theta}$, with $\theta$ representing the set of parameters that configure the digital encoder.
The optical encoder linearly projects the image to an intermediate vector $\textit{J}\in\mathbb{R}^{d}$ representing the actual measurements of the signal, where $d$ is 100 to 1000 times smaller than the pixel count $mn$ in the original image. This substantial compression ratio is achieved by optimizing the projection alongside other components during the image reconstruction process. Instead of merely matching pixel values, the process targets the generative model’s latent representation L of the original signal. By focusing on this compressed space, the essential features of the data distribution within a particular domain are captured, effectively eliminating redundancies inherent in the data.
However, due to the complex non-linear operations intrinsic to generative models and the impracticality of realizing these non-linearities optically, an additional digital encoder $\mathcal{D}_{\theta}$ is necessary to align J with the entire latent space of interest. This alignment is achieved through a deep neural network (DNN), which enhances the representation by effectively addressing the non-linearities.
The digital encoder $\mathcal{D}_{\theta}$ (see Fig.1d) expands J to $\textbf{L}\in\mathbb{R}^{512\times 18}$, aligning it with the chosen generative model’s dimensionality. In our work, StyleGAN2 [21] is employed since it performs robustly within specific domains, and its mechanisms for inversion to latent space are well-studied [23].
The digital encoder employs a multi-level network inspired by advancements in the inversion domain [24]. This approach leverages the detailed structural levels of StyleGAN2 [21]. In this framework, the optically encoded vector J passes through a series of linear layers and attention mechanisms, each followed by activation functions. The depth of these stacks varies to match the required level of detail. For the pre-trained StyleGAN2 on human faces, there are 18 levels corresponding to different resolutions. The number of stacks increases progressively from coarse to fine, aligning with low to high resolutions, respectively.
After acquiring all detail levels, we obtain $\textbf{L}\in\mathbb{R}^{512\times 18}$. A mixer block is then used to learn a weighted mixture of information across these levels, which is subsequently fed into the generative model $\mathcal{G}$ to output the image I’
@@
A pivotal aspect of the LSI framework is the identification of the optimal latent space to serve as ground-truth during the optimization process. To accomplish this, we employ a pre-trained encoder, $\mathcal{E}$, following the method proposed by [24] for accessing the StyleGAN2 latent space. This encoder provides an approximate ground-truth latent representation, $\mathcal{E}(\textit{I})$, which serves as the target for optimizing each entry of the O matrix and the $\mathcal{D}_{\theta}$. The optimization is driven by the loss function $\mathcal{L}_{lat}$
@@
$\begin{matrix}\mathcal{L}_{lat}=\left\|\mathcal{D}_{\theta}(\textbf{O}\cdot\textit{I}))-\mathcal{E}(\textit{I})\right\|_{1}\end{matrix}$ (3)
@@
Making the encoding L close to $\mathcal{E}(\textit{I})$ guarantees that it is inside the StyleGAN average space and diminishes hallucination caused by only matching an $\ell_{2}$ norm at the pixel image level. To further enhance facial fidelity, we incorporate additional loss functions. Detailed information on these loss functions can be found in the supplementary material.
Refer to caption a b c d
Figure 2: (a) Simulated results of our framework compared to the ground truth. The process involves optically encoding input images using learned masks. The resulting intermediate latent space vector is visualized through the “sea urchin” diagram, where each spike represents one measured variable. (b) Compression efficiency of the LSI framework demonstrated through side-by-side comparisons. The left side of the panel shows the ground truth, while the right side illustrates the progression from 512 to 64 measurements, representing compression rates from approximately $0.8\%$ to $0.1\%$ of the total pixel count. (c) Visual comparison across different imaging strategies: the input image, LSI, LSI with Variational Autoencoder (LSI-VAE), and a representative method from Compressive Sensing (CS) employing Fourier Single-Pixel Imaging (FSI). The FSI image is reconstructed using a deep learning algorithm from the AuSamNet [25] framework for fixed circular mask. This panel highlights the effectiveness of our imaging strategy in retrieving detailed features. (d) Quantitative comparison of identity retrieval capabilities between LSI and CS, using is the dot product between features extracted from CurricularFace [26], a specialized model for face recognition. This plot shows that LSI excels in high-compression settings but saturates at lower ratios, whereas CS exhibits poorer performance at extreme compression levels but improves substantially as the number of measurements increases.
Fig. 2a showcases the identity preservation capabilities of our model reconstructing images at a resolution of $256\times 256$ with $256$ measurements. Latent Space Imaging is able to retrieve eye colors, expression, shape, skin tone with less than $0.4\%$ of the original pixel count. This is achieved by designing O to retrieve specific features from the domain that we are targeting, thus pushing it to focus on more important characteristics on facial identification as well as the landmarks present in the underlying distribution learned by the generative model. This property, together with compressing the target space to be flat, drives the reconstruction process to not only interpolate under-sampled information but to produce an output close to the original detailed face.
Direct learning in the latent space is also an efficient method for handling color images, compared with traditional compressive sensing designs. In this approach, color is already implicitly represented in the latent space and does not require explicit modeling. O integrates the signal over the wavelengths to generate our measurement, which in simulation means to sum over the RGB channels. Typically, explicit modeling would increase the number of measurements necessary for reconstruction or necessitate some trade-off between spatial and spectral resolution. Latent Space Imaging (LSI) effectively bypasses these complexities, enabling going to further compressive levels utilizing only $512,256,128$, and $64$ measurements. Fig. 2b provides an example of detailed presence even in such critical situations. This contrasts with traditional compressive sensing strategies, where limited measurements lead to reconstruction algorithms over-smoothing the images, resulting in the loss of critical identity features such as the shape of eyes, beards, and other distinguishing landmarks. Fig. 2c provides visual evidence of this phenomenon, while Fig. 2d quantitatively demonstrates the differences. In this analysis, we compare Latent Space Imaging (LSI) against two other methods: LSI with a Variational Autoencoder (LSI-VAE) and Fourier Single-Pixel Imaging (FSI) [18]. For the FSI, we employ fixed circular masks coupled with a deep learning model for signal reconstruction, adhering to the procedures set forth in Huang et al.[25]. Detailed methodologies and additional comparisons are elaborated in supplementary material.
2.2 Physical Setup
Realizing the LSI framework involves implementing the optical encoder hardware along with the digital encoder software module. Among a range of different alternatives, we opt for the single-pixel camera approach as the hardware framework for our prototype. In this context, O spatially modulates the incoming image I through matrix multiplication (denoted by $\cdot$ in Equation 1). Each row of the optical encoder corresponds to a learned mask, and each column represents one measurement.
Utilizing the straight-through estimator (STE) [27], we enable the gradients to update the optical encoder, tailoring it for specific feature extraction while ensuring its physical viability. This involves maintaining O’s entries as positive, bounded between [0, 1], and quantized. To enhance the system’s light efficiency and control over subtle changes, an energy efficiency loss is strategically added during the later stages of training. This key adjustment finely tunes the encoding process, ensuring each pattern is distinctly counted, thereby boosting the system’s capability to discriminate between different patterns effectively. As demonstrated in Fig.3a, with the energy loss applied, the histogram showcases a broad spectrum of occupancies, reflecting a high diversity of patterns. In contrast, without this modification, the patterns are more uniform in their occupancy levels, exhibiting less variability. Additionally, Fig.3b illustrates that the masks projected onto the DMD following this loss application result in patterns that strikingly resemble faces, highlighting the system’s focus on facial landmarks for specific domain encoding. This enhancement not only augments detection capabilities but also prevents minor intensity changes in the projected image from going unnoticed.

### v2

@@ Abstract
Digital imaging systems have traditionally relied on brute-force measurement and processing of pixels arranged on regular grids. In contrast, the human visual system performs significant data reduction from the large number of photoreceptors to the optic nerve, effectively encoding visual information into a low-bandwidth latent space representation optimized for brain processing. Inspired by this, we propose a similar approach to advance artificial vision systems.
@@
Latent Space Imaging introduces a new paradigm that combines optics and software to encode image information directly into the semantically rich latent space of a generative model. This approach substantially reduces bandwidth and memory demands during image capture and enables a range of downstream tasks focused on the latent space.
@@
We validate this principle through an initial hardware prototype based on a single-pixel camera. By implementing an amplitude modulation scheme that encodes into the generative model’s latent space, we achieve compression ratios ranging from 1:100 to 1:1000 during imaging, and up to 1:16384 for downstream applications. This approach leverages the model’s intrinsic linear boundaries, demonstrating the potential of latent space imaging for highly efficient imaging hardware, adaptable future applications in high-speed imaging, and task-specific cameras with significantly reduced hardware complexity.
1 Introduction
@@
Conventional imaging systems are based on brute-force sampling of scene information on regular grids. In contrast, biological vision systems operate very differently: for example the human visual system has around 120 million rods and 7 million cones [44], but the information from these photoreceptors is distilled by retinal processing into neural signals that are transmitted over the optic nerve with only 0.7-1.7 million axions [25]. Therefore, our vision employs compression to encode image information into a latent space compatible with the neural architecture of the human brain.
@@
Here we propose Latent Space Imaging (LSI), a new paradigm that aims to bring this concept to the realm of cameras and artificial vision systems. Specifically, we propose to utilize the combination of linear optical encoding for drastic data reduction, paired with minimal non-linear computational processing to directly encode image information into the latent space of a generative model. This approach extremely reduces the amount of sensor data required for downstream tasks such as face segmentation, facial landmark detection, or image reconstruction.
To showcase this framework, we focus here on the rich latent space of a generative model in the form of a StyleGANXL [28, 29, 30, 53] trained on human faces. StyleGAN models are renowned for their ability to synthesize highly realistic face images, while supporting editing and image-to-image translation applications. Their semantically rich, disentangled, well-structured, and compact latent space makes them an ideal architecture for LSI. This structured representation facilitates key downstream applications, including landmark detection, facial feature segmentation, and the extraction of facial attributes such as age, gender, smile/no smile, and beard/no beard. Furthermore, we demonstrate its effectiveness in the particularly challenging task of reconstructing full human face images from a very small number of samples.
While there are a number of possible ways to optically implement the LSI paradigm, our first prototype utilizes the single-pixel imaging (SPI) concept [15, 31, 59, 77]. A discussion of alternative implementations [68, 6, 40, 67] can be found in the Supplement.
In summary, we make the following contributions:
  • the introduction of Latent Space Imaging as a new paradigm for very low bandwidth image capture;
  • a proof-of-concept hardware prototype for LSI utilizing the SPI framework, featuring a programmable optical system for easy and extensive experimentation,
  • an initial latent space for human faces based on a StyleGAN’s family of generative model; and
  • the demonstration of a range of downstream applications on this latent space with real hardware experimentation.
We also refer the readers to the Supplement for a detailed discussion of hardware and software choices for the LSI paradigm. Our implementation is available at https://github.com/vccimaging/latent-imaging.
2 Related Work
2.1 Optics and Algorithm Co-design
@@
Latent Space Imaging builds and improves upon the recent trend for co-designing both optics and reconstruction algorithms. This joint optimization has yielded favorable outcomes in computational imaging and downstream computer vision tasks. These include color image restoration [4], microscopy [46, 35, 55], monocular depth imaging [5, 18, 19, 69], super-resolution, extended depth of field [57, 61], time-of-flight imaging [10], and high-dynamic range imaging [60], among others. The enhancements observed are attributed to the ability of such systems to optically encode relevant scene information into sensor data, with the computational post-processing acting as a decoder. The crucial innovation in LSI is that the code represents the semantically expressive yet efficiently encoded latent space of a generative model, with that model acting as the decoder module.
2.2 Compressive Sensing and Minimal Cameras
Our approach aligns with the principles of Compressive Sensing (CS) [14], which harnesses sparsity and structural priors to recover signals from minimal samples. Traditional CS methods emphasize optimization-based reconstruction, while recent advancements integrate deep learning, combining sparsity-driven priors with neural network capabilities. This first group of methods has led to diverse architectures, including CNNs, transformers, and deep unfolding models [74, 45, 58, 64, 49, 66, 65], bridging the gap between classical CS and data-driven approaches. Further, generative priors, forming a second group, offer a robust alternative by leveraging latent spaces to enhance image quality and structure [3, 26, 24, 34, 33], showing the potential of learned priors to capture dependencies within compressed data. A third group advances these concepts in real-world applications through single-pixel imaging [15, 31, 59, 77, 21, 22, 52, 73], using quantized patterns, binary random masks, Fourier domain methods, and co-optimized mask-reconstruction pipelines. Lastly, a fourth group [37, 48] presents minimalist cameras with optimized freeform pixels for task-specific applications like workspace monitoring.
Latent Space Imaging stands apart from both the first and second groups by emphasizing realistic physical deployment, requiring attention to numerical constraints, quantization, light efficiency, and similar practical elements. Unlike the second group, which often employs iterative optimization, our approach reconstructs images in a single forward pass. Additionally, we distinguish our work from these groups by targeting latent space reconstruction itself. Our framework uniquely leverages once-optimized masks for multi-model purposes, avoiding the conventional use of latent space solely as an optimization prior and without optimizing the masks/pixels for multiple tasks. Finally, we set ourselves apart from the fourth group design by utilizing a latent space to enable fine-grained tasks, such as face identification, with a smaller number of freeform samples than traditional cameras.
3 Latent Space Imaging
@@
Traditional imaging captures a regular $m\times n$ grid of measurements, while compression strategies typically use a collection of random linear codes to leverage the information redundancy inherent in real images in combination with a reconstruction software module. In contrast, we demonstrate that meaningful information can be significantly more compressed in the latent space of generative models while providing sufficient information for downstream processing tasks.
@@
As depicted by Fig. LABEL:fig:teaser, our objective is to learn an optically encoded mapping
@@
where the input image $\textit{I}\in\mathbb{R}^{mn}$ is mapped to a tensor $\textbf{L}\in\mathbb{R}^{k\times l}$ representing the latent space of a generative model. This mapping is achieved through the combination of a co-designed (linear) optical encoder $\textbf{O}\in\mathbb{R}^{d\times mn}$ and a digital (nonlinear) encoder $\mathcal{D}_{\theta}$, with $\theta$ representing the set of parameters that configure the digital encoder. Where $d$ is the dimension of the compressed measurement, $l$ is the number levels from coarse to fine inside the generative models as well as in the latent space, and $k$ represents latent dimension.
[…]
However, due to the complex non-linear operations intrinsic to generative models and the impracticality of realizing these non-linearities optically, an additional digital encoder $\mathcal{D}_{\theta}$ is necessary to align J with the entire latent space of interest. This alignment is achieved through a deep neural network (DNN), that enhances the representation by effectively addressing the non-linearities. The digital encoder $\mathcal{D}_{\theta}$ expands J to $\textbf{L}\in\mathbb{R}^{512\times 18}$ (see Fig. LABEL:fig:teaser), aligning it with the dimensions of the chosen generative model. In our work, StyleGANXL [53] is employed since it performs robustly within specific domains, and its mechanisms for inversion to latent space are well-studied [70].
Specifically, the digital encoder employs a multi-level network inspired by advancements in the inversion domain [51]. This approach leverages the detailed structural levels of StyleGANXL. In this framework, the optically encoded vector J passes through a series of linear layers and attention mechanisms, each followed by activation functions. The depth of these stacks varies to match the required level of detail. For the pre-trained StyleGANXL on human faces, there are 18 levels corresponding to different resolutions. The number of stacks increases progressively from coarse to fine, aligning with low to high resolutions, respectively.
After acquiring all detail levels, we obtain $\textbf{L}\in\mathbb{R}^{512\times 18}$. A mixer block, inspired by [41], is then used to learn a weighted mixture of information across these levels, which is subsequently fed into the generative model $\mathcal{G}$ to output the image I’
@@
A pivotal aspect of the LSI framework is the identification of the optimal latent space to serve as ground-truth during the optimization process. To accomplish this, we trained an encoder, $\mathcal{E}$, following the method proposed by [51] for accessing the StyleGANXL latent space. This encoder provides an approximate ground-truth latent representation, $\mathcal{E}(\textit{I})$, which serves as the target for optimizing each entry of the O matrix and the $\mathcal{D}_{\theta}$. The optimization is driven by the loss function $\mathcal{L}_{lat}$
@@
$\begin{matrix}\mathcal{L}_{lat}=\left\|\mathcal{D}_{\theta}(\textbf{O}\cdot\textit{I}))-\mathcal{E}(\textit{I})\right\|_{1}.\end{matrix}$ (3)
@@
Making the encoding L close to $\mathcal{E}(\textit{I})$ guarantees that it is inside the generative model average space and diminishes hallucination caused by only matching an $\ell_{2}$ norm at the pixel level. To further enhance facial fidelity, we incorporate additional terms in the loss function. Detailed information on these loss functions can be found in the Supplement.
4 Linear Learning for Downstream Tasks
In the LSI pipeline, we use a hybrid encoding setting to generate L, enabling high-level vision tasks with minimal training and a very limited number of measurements. Previous research has examined the semantics of GAN models, uncovering latent space linear boundaries that distinguish different image attributes [71, 11, 56, 1]. In particular, [71] demonstrated that linear transformations can efficiently extract semantic information from feature maps without the need for image reconstruction or specialized encoders for various tasks.
Our approach involves performing three different high-level facial analysis tasks using simple linear transformations from a single latent space L. Unlike prior studies, we push compression limits by assessing the minimum measurements needed to derive meaningful facial semantics through linear projections $P$ from the same latent space. Once O and $\mathcal{D}_{\theta}$ are optimized, they are kept frozen and different $P$ are optimized for downstream applications such as:
Attribute Classification predicts facial attributes (e.g., age, beard, smile, gender) using a linear feature mapper $P_{A}$, which is implemented as a fully connected layer that projects $\textbf{L}\in\mathbb{R}^{512\times 18}$ to $\mathbb{R}^{40}$, representing 40 distinct facial characteristics.
Face Segmentation provides pixel-wise segmentation across facial regions (e.g., nose, eyes, face, mouth). To accurately capture spatial structure, we employ multi-level feature extraction from the generative model, combining these features through a linear projection block, $P_{S}$, which utilizes bilinear interpolation alongside convolutional layers. The generative model $\mathcal{G}$ offers feature maps at various resolutions, spanning from coarse ($36\times 36$) to fine ($256\times 256$). We extract features across six distinct levels, representing a comprehensive range of resolutions. This multi-level feature integration allows for precise semantic pixel assignment, effectively addressing the spatial limitations of raw L features, which alone lack the fine-grained detail essential for high-accuracy segmentation.
Landmark Detection yields the coordinates of key facial points within an image. Unlike segmentation, where multiple features are extracted, we focus here on a single coarse feature map from the generative model to capture spatial structures. This coarse feature is denoted by $\in\mathbb{R}^{1024\times 36\times 36}$ and is modulated by the latent space, L itself. A convolutional layer, combined with bilinear interpolation, projects and downsamples this feature to $\mathbb{R}^{68\times 8\times 8}$. Finally, a linear layer reduces the spatial dimensions, representing the 68 key points as 2D coordinates.
5 Hardware Prototype
Figure 1: We illustrate one possible implementation of the Latent Space Imaging technique using a single-pixel framework. An objective lens focuses the image onto the Digital Micromirror Device (DMD), which is responsible for implementing the learned mask to spatially modulate the incoming signal. This is followed by a relay lens, which focuses the modulated signal onto a photodiode (SPD) responsible for integrating the signal. Utilizing time multiplexing, we can retrieve the necessary measurements.
Realizing the LSI framework (as detailed in Fig. 1) involves implementing the optical encoder hardware along with the digital encoder software module. Among a range of different alternatives to realize this linear layer in real-world, we opt for the single-pixel camera approach as the hardware framework for our prototype, since it allows for programmable adjustments of the compression ratio for experimentation. The Supplement discusses alternative options for physical realization of the LSI concept.